# CaRiNG: Learning Temporal Causal Representation under Non-Invertible Generation Process

### v1

@@ Abstract
Identifying the underlying time-delayed latent causal processes in sequential data is vital for grasping temporal dynamics and making downstream reasoning. While some recent methods can robustly identify these latent causal variables, they rely on strict assumptions about the invertible generation process from latent variables to observed data. However, these assumptions are often hard to satisfy in real-world applications containing information loss. For instance, the visual perception process translates a 3D space into 2D images, or the phenomenon of persistence of vision incorporates historical data into current perceptions. To address this challenge, we establish an identifiability theory that allows for the recovery of independent latent components even when they come from a nonlinear and non-invertible mix. Using this theory as a foundation, we propose a principled approach, CaRiNG, to learn the Causal Representation of Non-invertible Generative temporal data with identifiability guarantees. Specifically, we utilize temporal context to recover lost latent information and apply the conditions in our theory to guide the training process. Through experiments conducted on synthetic datasets, we validate that our CaRiNG method reliably identifies the causal process, even when the generation process is non-invertible. Moreover, we demonstrate that our approach considerably improves temporal understanding and reasoning in practical applications.
Arxiv
@@ 1 Introduction
However, these nonlinear ICA-based methods usually assume that the mixing function (the generation process from sources to observations) is invertible, which may be difficult to satisfy in real-world scenarios, such as the 3D to 2D projection in the visual process. As shown in Figure 1 (a) and (b), we provide two intuitive instances of the real videos to illustrate how the non-invertibility happens. In (a), when object occlusions occur, information from the obstructed object is lost in the generation process of the current time step, which causes non-invertibility. In (b), the persistence of vision introduces non-invertibility, since the mixing process of the current time step utilizes the history information. We further found that the violation of this inevitability assumption may cause the nonlinear ICA method to yield poor identification performance. In part (c) of Figure 1, we demonstrate that TDRL, one of the typical nonlinear ICA-based methods making the invertibility assumption, markedly degrades its performance in identifying the latent variables with increasing non-invertibility. It motivates us to extend the current nonlinear ICA methods to consider non-invertible mixing function.
[…]
@@
In this paper, to tackle the challenges above, we propose to leverage the temporal context for retrieving missing information caused by the non-invertible mixing function, mirroring the intuitive mechanisms of human perception. For instance, when we encounter an object with occlusion, our natural inclination is to draw from historical data to reconstruct the obscured portion. We demonstrate that, even when the generation process is non-invertible, the derived latent causal representation remains identifiable if the latent variables can be expressed as an arbitrary function combining the current observation with its history. Built upon this identification theorem, we introduce a principled approach, named CaRiNG, that learns the function to integrate historical data to compensate for the latent information lost due to non-invertibility. This approach extends the Sequential Variational Autoencoder (Sequential VAE) (Chung et al., 2015) with two distinct modifications. Firstly, it incorporates history (or context) information directly into the encoder. Specifically, we transform step-to-step mapping (from current observation to the current latent variable) into sequence-to-step mapping (from current observation and temporal context to the current latent variable). Secondly, a specialized prior module is introduced to determine the prior distribution of latent variables using the normalizing flow (Dinh et al., 2016), ensuring the imposition of an independent noise condition. We evaluate our method using both synthetic and real-world data. Using synthetic data, we design datasets with a non-invertible mixing function to gauge identifiability. For real-world applications, CaRiNG is deployed in a traffic accident reasoning task, a scenario in which the intricate traffic dynamics introduce considerable non-invertibility. Experimental outcomes reveal that our method significantly outperforms other temporal representation learning methods for identifying causal representations amid non-invertible generation processes. Furthermore, this causal representation has proven instrumental in enhancing video reasoning tasks.
@@
  • To the best of our understanding, this paper presents the first identifiability theorem that accommodates a non-invertible generation process, which complements the existing body of the nonlinear ICA theory, and enables more complex applications.
@@
  • We conduct the experiments on both synthetic and real-world datasets. The evaluation results demonstrate the CaRiNG’s effectiveness for learning the identifiable latent causal representation, leading to enhancements in video reasoning tasks.
@@ 2.1 Non-invertible Temporal Generative Process
Consider observing $n$-dimensional time series data with $T$ discrete time steps , represented as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$. Each observation $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is generated from a nonlinear mixing function $\mathbf{g}$ that maps $r+1$ adjacent latent variables $\mathbf{z}_{t:t-r}$ to $\mathbf{x}_{t}$, where $\mathbf{z}_{t:t-r}$ refers to $\{\mathbf{z}_{t},\mathbf{z}_{t-1},\cdots,\mathbf{z}_{t-r}\}$. We have $\mathbf{z}_{t}\in\mathbb{R}^{n}$. For every $i\in{1,\dots,n}$, the variable $z_{it}$ of $\mathbf{z}_{t}$ is derived from a stationary, non-parametric time-delayed causal relation:
@@
$\underbrace{z_{it}=f_{i}\left(\{z_{j,t-\tau}|z_{j,t-\tau}\in\mathbf{Pa}(z_{it})\},\epsilon_{it}\right)}_{\text{Stationary non-parametric transition}}.$
@@ 2.2 Identification of the Latent Causal Processes
Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ be a sequence of observed variables generated by the true temporally causal latent processes specified by $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ given in Eq 1. A learned generative model $(\hat{f}_{i},\hat{p}({\epsilon_{i}}),\hat{\mathbf{g}})$ is observational equivalent to $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ if the model distribution $p_{\hat{f}_{i},\hat{p}_{\epsilon},\hat{\mathbf{g}}}(\mathbf{x}_{1:T})$ matches the data distribution $p_{f_{i},p_{\epsilon},\mathbf{g}}(\mathbf{x}_{1:T})$ for any value of $\mathbf{x}_{t}$. We say latent causal processes are identifiable if observational equivalence can lead to a version of latent variable $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ up to permutation $\pi$ and component-wise invertible transformation $T$:
@@
$\Rightarrow$ $\hat{\mathbf{m}}(\mathbf{x}_{t:t-\mu})\!=\!(T\!\circ\!\pi\!\circ\!\mathbf{m})(\mathbf{x}_{t:t-\mu}),\forall\mathbf{x}_{t:t-\mu}\!\in\!\mathbb{R}^{\mu+1}.$
Different from the existing literature, we involve $\mathbf{m}$ in the above definition, since it does so implicitly as a property of the mixing function $\mathbf{g}$, although it does not explicitly participate in the generation process. Furthermore, the identifiability of $\mathbf{g}$ is different. In previous nonlinear ICA methods (Yao et al., 2022a; Hyvarinen & Morioka, 2017), the mixing function $\mathbf{g}$ is identifiable. However, in our case, we cannot find the identifiable mixing function since the information loss is caused by non-invertibility. Instead, we can obtain a component-wise transformation of a permuted version of latent variables $\mathbf{\hat{z}}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$. The latent causal relations are also identifiable, up to a permutation $\pi$ and component-wise invertible transformation $T$, i.e., $\mathbf{\hat{f}}=T\circ\pi\circ\mathbf{f}$, once $\mathbf{z}_{t}$ is identifiable. Because in the time-delayed causally sufficient system, the conditional independence relations fully characterize time-delayed causal relations when we assume no latent causal confounders in the (latent) causal processes.
Figure 2: An intuitive illustration of a moving ball with a visual persistence effect. The ball is moving along a descending curve. The observation in a particular time step is composed of multiple residual images of the ball.
@@ 2.3 Illustrations of the Problem Setup
Illustration of time-delayed temporal relations. Here, we assume that there are only time-delayed temporal relations in the time series system. In other words, any instantaneous relations will not fall into the discussion. Generally speaking, a group of objects which have instantaneous relations with each other would be treated as one single variable. For example, within a video sequence, a ball in motion may be conceptualized as a cluster of pixels that move consistently and simultaneously (instantaneous relations). This pattern can help distinguish the ball from the others, which potentially provides a principle to extract concepts from time series data like video, motion sequence, etc.
@@ 3 Identifiability Theory
In this section, we demonstrate that, given certain mild conditions, the learned causal representation $\mathbf{z}_{t}$ is identifiable up to permutation and a component-wise transformation. This holds even if the mixing function $\mathbf{g}$ is non-invertible. Firstly, we present the identifiability results when faced with a non-invertible mixing function and stationary transitions. Subsequently, we address the gap between permutation-scaling Jacobian and identifiability. Lastly, by leveraging side information such as the domain index and label, we illustrate how identifiability can be achieved even in a non-stationary context. The exhaustive proofs are available in Appendix A.
@@ 3.1 Identifiability under Non-Invertible Generative Process
where a function $\mathbf{m}$ satisfying $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ exists. When taking $r=0$, the time delay is present only in transitions and is absent in the generation process. Taking $r>0$ leads us to a more intricate scenario, where the mixing function encompasses not just the latent causal variables of the current time step, but also the information of previous steps, termed the Time-delayed Mixing Process. Such a scenario is compelling, acknowledging that the mixing process can be influenced by time-delayed effects. To illustrate, human visual perception provides a fitting example: the phenomenon known as the persistence of vision reveals that humans retain impressions of a visual stimulus even after its cessation (Coltheart, 1980). The extensions for any time lag $\tau$ will be discussed in Appendix A.2.
@@
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ whose probability density function is smooth and positive, suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which subject to observational equivalence, i.e.,
@@
[…]
    for each value of $\mathbf{z}_{t}$, $\mathbf{v}_{1t},\mathring{\mathbf{v}}_{1t},\mathbf{v}_{2t},\mathring{\mathbf{v}}_{2t},...,\mathbf{v}_{nt},\mathring{\mathbf{v}}_{nt}\in\mathbf{R}^{2n}$, as $2n$ vector functions in $z_{1,t-1}$, $z_{2,t-1}$, …, $z_{n,t-1}$, are linearly independent
@@
The proof of Theorem 1 can be found in Appendix A.1. It is inspired from (Yao et al., 2022a), which follows the line of (Hyvarinen et al., 2019).
Besides, the nonstationary transition can also help to improve the identifiability of CaRiNG. As shown in the sufficiency assumption in Theorem 1, the identifiability relies on the sufficient changes of the conditional distribution $p(z_{kt}|\mathbf{z}_{t-1:t-r-1})$. When the distribution of the noise term varies between different domains, the domain index can serve as an auxiliary variable to improve this sufficiency since both domain dynamics and history variables can provide changes. A further discussion is provided in the Appendix A.4.
@@ 3.2 Continuity for Permutation Invariance
Following Definition 1, if $\pi$ is a fixed permutation and $T$ is a component-wise invertible transition which may vary across different time steps. We call this identifiability under Permutation Invariance.
Let us further consider a more general scenario, with $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$, i.e., the lebesgue measure of $\mathbf{z}_{t}$ does not have to be non-zero everywhere in $\mathbb{R}^{n}$ To establish identifiability, numerous existing nonlinear ICA-based methods (Yao et al., 2022b, a; Hyvarinen et al., 2019; Hälvä et al., 2021) utilize the Jacobian matrix, denoted by $\mathbf{H}=\frac{\partial\mathbf{z}}{\partial\hat{\mathbf{z}}}$, which captures the relationship between ground truth and estimated latent variables. These methods propose that the learned latent variables are identifiable if $\mathbf{H}_{ij}\cdot\mathbf{H}_{ik}=0$ for $j\not=k$ (with only a single non-zero element in each row or column). $\mathbf{H}$ corresponds to the Jacobian matrix of the function $\mathbf{h}\triangleq\mathbf{m}\circ\hat{\mathbf{g}}$ in our scenario (or $\mathbf{g}^{-1}\circ\hat{\mathbf{g}}$ for the general scenario). However, it is crucial to highlight an often overlooked shortcoming: this condition alone is insufficient to establish identifiability when dealing with non-linear generation processes. Concurrently to our work, (Lachapelle et al., 2023) also arrived at the difference between local and global disentanglement, and achieved the global disentanglement under the additive decoding case. Alternatively, we demonstrate the identifiability under the permutation invariance and focus on a more general case without the block-specific decoder assumptions. While in linear ICA, given that the Jacobian remains constant, this condition indeed equates to identifiability. Yet, in nonlinear ICA, the Jacobian matrix, being a function of $\mathbf{\hat{z}}$, can vary with different $\mathbf{\hat{z}}$ values, potentially rendering the mapping unpredictable. A comprehensive discussion is available in Appendix A.3.
@@
A function $\mathbf{z}=\mathbf{h}(\mathbf{\hat{z}},\mathbf{c})$, where $\mathbf{z},\mathbf{\hat{z}}\in\mathbb{R}^{n}$ and $\mathbf{z}\in\mathbb{R}^{m}$, is partially invertible, if and only if for any given $\mathbf{c}$, the unfixed part $\mathbf{h}_{\mathbf{c}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is always invertible.
@@
For a series of observations $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$, suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which subject to observational equivalence, i.e.,

### v2

@@ Abstract
Identifying the underlying time-delayed latent causal processes in sequential data is vital for grasping temporal dynamics and making downstream reasoning. While some recent methods can robustly identify these latent causal variables, they rely on strict assumptions about the invertible generation process from latent variables to observed data. However, these assumptions are often hard to satisfy in real-world applications containing information loss. For instance, the visual perception process translates a 3D space into 2D images, or the phenomenon of persistence of vision incorporates historical data into current perceptions. To address this challenge, we establish an identifiability theory that allows for the recovery of independent latent components even when they come from a nonlinear and non-invertible mix. Using this theory as a foundation, we propose a principled approach, CaRiNG, to learn the Causal Representation of Non-invertible Generative temporal data with identifiability guarantees. Specifically, we utilize temporal context to recover lost latent information and apply the conditions in our theory to guide the training process. Through experiments conducted on synthetic datasets, we validate that our CaRiNG method reliably identifies the causal process, even when the generation process is non-invertible. Moreover, we demonstrate that our approach considerably improves temporal understanding and reasoning in practical applications. Code can be accessed through https://github.com/sanshuiii/CaRiNG.
Machine Learning, ICML
@@ 1 Introduction
However, these nonlinear ICA-based methods usually assume that the mixing function (the generation process from sources to observations) is invertible, which may be difficult to satisfy in real-world scenarios, such as the 3D to 2D projection in the visual process. As shown in Figure 1 (a) and (b), we provide two intuitive instances of the real videos to illustrate how the non-invertibility happens. In (a), when object occlusions occur, information from the obstructed object is lost in the generation process of the current time step, which causes non-invertibility. In (b), the persistence of vision introduces non-invertibility, since the mixing process of the current time step utilizes the history information. We further found that the violation of this invertibility assumption may cause the nonlinear ICA method to yield poor identification performance. In part (c) of Figure 1, we demonstrate that TDRL, one of the typical nonlinear ICA-based methods making the invertibility assumption, markedly degrades its performance in identifying the latent variables with increasing non-invertibility. It motivates us to extend the current nonlinear ICA methods to consider non-invertible mixing function.
Figure 1: Motivations of the non-invertible generation process. (a) The occlusions raise the non-invertibility since the measured observation cannot cover the obstructed objects. (b) The vision persistence, shown with the high-speed movement of a crashing car, describes the generation process that jointly involves the current state and previous, and causes the non-invertibility. (c) The identifiability of conventional methods, such as TDRL (Yao et al., 2022a) (blue), drops drastically with the increase of non-invertibility, while the identifiability of our method (marked in orange) still holds. The levels of non-invertibility are defined by removing $0,1/3$, and $2/3$ dimensions of $\bf{z}_{t}$ when generating $\bf{x}_{t}$. For example, when the dimension of $\bf{z}_{t}$ is 6, $2/3$ non-invertibility means that we remove 4 variables of $\bf{z}_{t}$ and use only 2 variables to generate $\bf{x}_{t}$.
@@
In this paper, to tackle the challenges above, we propose to leverage the temporal context for retrieving missing information caused by the non-invertible mixing function, mirroring the intuitive mechanisms of human perception. For instance, when we encounter an object with occlusion, our natural inclination is to draw from historical data to reconstruct the obscured portion. We demonstrate that, even when the generation process is non-invertible, the derived latent causal representation remains identifiable if the latent variables can be expressed as an arbitrary function combining the current observation with its history. Built upon this identification theorem, we introduce a principled approach, named CaRiNG, that learns the function to integrate historical data to compensate for the latent information lost due to non-invertibility. This approach extends the Sequential Variational Autoencoder (Sequential VAE (Chung et al., 2015; Li & Mandt, 2018)) with two distinct modifications. Firstly, it incorporates history (or context) information directly into the encoder. Specifically, we transform step-to-step mapping (from current observation to the current latent variable) into sequence-to-step mapping (from current observation and temporal context to the current latent variable). Secondly, a specialized prior module is introduced to determine the prior distribution of latent variables using the normalizing flow (Dinh et al., 2016), ensuring the imposition of an independent noise condition. We evaluate our method using both synthetic and real-world data. Using synthetic data, we design datasets with a non-invertible mixing function to measure identifiability. For real-world applications, CaRiNG is deployed in a traffic accident reasoning task, a scenario in which the intricate traffic dynamics introduce considerable non-invertibility. Experimental outcomes reveal that our method significantly outperforms other temporal representation learning methods for identifying causal representations amid non-invertible generation processes. Furthermore, this causal representation has proven instrumental in enhancing video reasoning tasks.
@@
  • To the best of our understanding, this paper presents the first identifiability theorem that accommodates a non-invertible generation process, which complements the existing body of the nonlinear ICA theory.
@@
  • Our evaluations across synthetic and real-world datasets demonstrate the CaRiNG’s effectiveness for learning the identifiable latent causal representation, leading to enhancements in video reasoning tasks.
@@ 2.1 Non-invertible Temporal Generative Process
Denote $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ as the observed $d$-dimensional time series data at $T$ discrete time steps. Each observation $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is generated from a nonlinear mixing function $\mathbf{g}$ that maps $r+1$ adjacent latent variables $\mathbf{z}_{t:t-r}$ to $\mathbf{x}_{t}$, where $\mathbf{z}_{t:t-r}$ refers to $\{\mathbf{z}_{t},\mathbf{z}_{t-1},\cdots,\mathbf{z}_{t-r}\}$. We have $\mathbf{z}_{t}\in\mathbb{R}^{n}$. For every $i\in{1,\dots,n}$, the variable $z_{it}$ of $\mathbf{z}_{t}$ is derived from a stationary, non-parametric time-delayed causal relation:
@@
$\underbrace{z_{it}={f_{i}\left(\{z_{j,t^{\prime}}|z_{j,t^{\prime}}\in\mathbf{Pa}(z_{it})\},\epsilon_{it}\right)}}_{\text{Stationary non-parametric transition}}.$
@@ 2.2 Identification of the Latent Causal Processes
Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ be a sequence of observed variables generated by the true temporally causal latent processes specified by $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ given in Eq 1. A learned generative model $(\hat{f}_{i},\hat{p}({\epsilon_{i}}),\hat{\mathbf{g}})$ is observational equivalent to $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ if the model distribution $p_{\hat{f}_{i},\hat{p}_{\epsilon},\hat{\mathbf{g}}}(\mathbf{x}_{1:T})$ matches the data distribution $p_{f_{i},p_{\epsilon},\mathbf{g}}(\mathbf{x}_{1:T})$ for any value of $\mathbf{x}_{t}$. We say latent causal processes are identifiable if observational equivalence can lead to a version of latent variable $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ up to permutation $\pi$ and component-wise invertible transformation $\mathcal{T}$:
@@
$\Rightarrow$ $\hat{\mathbf{m}}(\mathbf{x}_{t:t-\mu})\!=\!({\mathcal{T}}\!\circ\!\pi\!\circ\!\mathbf{m})(\mathbf{x}_{t:t-\mu}),\forall\mathbf{x}_{t:t-\mu}\!\in\!\mathbb{R}^{\mu+1}.$
Different from the existing literature, we involve $\mathbf{m}$ in the above definition since it serves implicitly as a property of the mixing function $\mathbf{g}$, although it does not explicitly participate in the generation process. Furthermore, the identifiability of $\mathbf{g}$ is different. In previous nonlinear ICA methods (Yao et al., 2022a; Hyvarinen & Morioka, 2017), the mixing function $\mathbf{g}$ is identifiable. However, in our case, we cannot find the identifiable mixing function since the information loss is caused by non-invertibility. Instead, we can obtain a component-wise transformation of a permuted version of latent variables $\mathbf{\hat{z}}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$. The latent causal relations are also identifiable, up to a permutation $\pi$ and component-wise invertible transformation $\mathcal{T}$, i.e., $\mathbf{\hat{f}}=\mathcal{T}\circ\pi\circ\mathbf{f}$, once $\mathbf{z}_{t}$ is identifiable. Because in the time-delayed causally sufficient system, the conditional independence relations fully characterize time-delayed causal relations when we assume no latent causal confounders in the (latent) causal processes.
Figure 2: An intuitive illustration of a moving football with a visual persistence effect. Considering the generating process $\mathbf{x}_{t}=\mathbf{g}(\mathbf{z}_{t:t-r})$, $\mathbf{x}_{t}$ denotes the observed football with motion blur, and $\mathbf{z}_{t}$ denotes the position and phase of the ball. Recovering the latent variables from a single observation will be difficult, which introduces non-invertibility.
@@ 2.3 Illustrations of the Problem Setup
Illustration of time-delayed temporal relations. Here, we assume that there are only time-delayed temporal relations in the time series system. In other words, any instantaneous relations will not fall into the discussion. Generally speaking, a group of objects that have instantaneous relations with each other would be treated as one single variable. For example, within a video sequence, a ball in motion may be conceptualized as a cluster of pixels that move consistently and simultaneously (instantaneous relations). This pattern can help distinguish the ball from the others, which potentially provides a principle to extract concepts from time series data like video, motion sequence, etc.
@@ 3 Identifiability Theory
In this section, we demonstrate that, given certain mild conditions, the learned causal representation $\mathbf{z}_{t}$ is identifiable up to permutation and a component-wise transformation. This holds even if the mixing function $\mathbf{g}$ is non-invertible. Firstly, we present the identifiability results when faced with a non-invertible mixing function and stationary transitions. Subsequently, we address the gap between permutation-scaling Jacobian and identifiability. Lastly, by leveraging side information such as the domain index and label, we illustrate how identifiability can be achieved even in a non-stationary context. The proofs are available in Appendix A1.
@@ 3.1 Identifiability under Non-Invertible Generative Process
where a function $\mathbf{m}$ satisfying $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ exists. When taking $r=0$, the time delay is present only in transitions and is absent in the generation process. Taking $r>0$ leads us to a more intricate scenario, where the mixing function encompasses not just the latent causal variables of the current time step, but also the information of previous steps, termed the Time-delayed Mixing Process. Such a scenario is compelling, acknowledging that the mixing process can be influenced by time-delayed effects. To illustrate, human visual perception provides a fitting example: the phenomenon known as the persistence of vision reveals that humans retain impressions of a visual stimulus even after its cessation (Coltheart, 1980). The extensions for any time lag $\tau$ will be discussed in Appendix A1.4.
@@
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{d}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$, suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which is subject to observational equivalence,
@@
  • (Smooth and Positive Density) the probability density function of latent variables is third-order differentiable and positive in $R^{n}$,
@@
    $\mathbf{v}_{lt}\triangleq\Big{(}\frac{\partial^{2}\eta_{1t}}{\partial z_{1t}\partial z_{l,t-r-1}},...,\frac{\partial^{2}\eta_{nt}}{\partial z_{nt}\partial z_{l,t-r-1}},$ (6)
    $\frac{\partial^{3}\eta_{1t}}{\partial z_{1t}^{2}\partial z_{l,t-r-1}},...,\frac{\partial^{3}\eta_{nt}}{\partial z_{nt}^{2}\partial z_{l,t-r-1}}\Big{)}^{\intercal},$
    for $l=1,2,\cdots,n$. For each value of $\mathbf{z}_{t}$, there exists $2n$ different values of $z_{l,t-r-1}$ such that the $2n$ vector functions $\mathbf{v}_{lt}\in\mathbf{R}^{2n}$ are linearly independent,
@@
The proof of Theorem 1 can be found in Appendix A1.1. It is inspired by (Yao et al., 2022a), which follows the line of (Hyvarinen et al., 2019).
Besides, the nonstationary transition can also help to improve the identifiability of CaRiNG. As shown in the sufficiency assumption in Theorem 1, the identifiability relies on the sufficient changes of the conditional distribution $p(z_{kt}|\mathbf{z}_{t-1:t-r-1})$. When the distribution of the noise term varies between different domains, the domain index can serve as an auxiliary variable to improve this sufficiency since both domain dynamics and history variables can provide changes. More discussions are in Appendix A1.6.
@@ 3.2 Continuity for Permutation Invariance
Following Definition  1, if $\pi$ is a fixed permutation and $\mathcal{T}$ is a component-wise invertible transformation which may vary across different time steps, we call this identifiability under Permutation Invariance.
Let us further consider a more general scenario, with $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$, i.e., the probability density of $\mathbf{z}_{t}$ does not have to be non-zero everywhere in $\mathbb{R}^{n}$ To establish identifiability, numerous existing nonlinear ICA-based methods (Yao et al., 2022b, a; Hyvarinen et al., 2019; Hälvä et al., 2021) utilize the Jacobian matrix, denoted by $\mathbf{H}=\frac{\partial\mathbf{z}}{\partial\hat{\mathbf{z}}}$, which captures the relationship between ground truth and estimated latent variables. These methods propose that the learned latent variables are identifiable if $\mathbf{H}_{ij}\cdot\mathbf{H}_{ik}=0$ for $j\not=k$ (with only a single non-zero element in each row or column). $\mathbf{H}$ corresponds to the Jacobian matrix of the function $\mathbf{h}\triangleq\mathbf{m}\circ\hat{\mathbf{g}}$ in our scenario (or $\mathbf{g}^{-1}\circ\hat{\mathbf{g}}$ for the general scenario). However, it is crucial to highlight an often overlooked shortcoming: this condition alone is insufficient to establish identifiability when dealing with non-linear generation processes. Concurrently to our work, (Lachapelle et al., 2023) also arrived at the difference between local and global disentanglement, and achieved the global disentanglement under the additive decoding case. Alternatively, we demonstrate the identifiability under the permutation invariance and focus on a more general case without the block-specific decoder assumptions. While in linear ICA, given that the Jacobian remains constant, this condition indeed equates to identifiability. Yet, in nonlinear ICA, the Jacobian matrix, being a function of $\mathbf{\hat{z}}$, can vary with different $\mathbf{\hat{z}}$ values, potentially rendering the mapping unpredictable. A comprehensive discussion is available in Appendix A1.5.
@@
A function $\mathbf{z}=\mathbf{h}(\mathbf{\hat{z}},\mathbf{c})$, where $\mathbf{z},\mathbf{\hat{z}}\in\mathbb{R}^{n}$ and $\mathbf{c}\in\mathbb{R}^{m}$, is partially invertible, if and only if for any given $\mathbf{c}$, the rest part $\mathbf{h}_{\mathbf{c}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is always invertible.
@@
For a series of observations $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$, suppose there exists function $\mathbf{\hat{g}},\mathbf{\hat{m}}$ which subject to observational equivalence, i.e.,